# New Paradigm of Adversarial Training: Breaking Inherent Trade-Off between Accuracy and Robustness via Dummy Classes

### v1

@@ 2.2.1 Preliminaries
In the context of clean training, for a classification task with $C\geq 2$ as the number of classes, given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1,...,n}$ with $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{1,...,C\}$ respectively denoting a natural sample and its supervised label, as well as $\mathbf{y}_{i}\in\mathbb{N}^{C}$ being the one-hot format of $y_{i}$, ERM aims to learn a classifier $h_{\bm{\theta}}:\mathbf{x}\rightarrow\mathbb{R}^{C}$, such that $h_{\bm{\theta}}(\mathbf{x}_{i})=\mathbf{q}(\mathbf{x}_{i},\bm{\theta})$ represents the output logits with each element $\mathbf{q}^{(k)}(\mathbf{x}_{i},\bm{\theta})$ corresponding to class $k$, to minimize the empirical risk $\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{D}}\,[\mathcal{L}(h_{\bm{\theta}}(\mathbf{x}),\mathbf{y})]$ through a classification loss $\mathcal{L}$ like cross-entropy (CE). Such a learning objective is formulated as:
@@
In contrast, AT augments conventional ERM to $\mathbb{E}_{\mathbf{x}^{\prime},\mathbf{y}\sim\mathcal{D}}\,[\max_{\,\mathbf{x}^{\prime}\in\mathcal{P}(\mathbf{x})}\mathcal{L}(h_{\bm{\theta}}(\mathbf{x}^{\prime}),\mathbf{y})]$, where $\mathbf{x}^{\prime}$ denotes the adversarial sample w.r.t. $\mathbf{x}$ and $\mathcal{P}(\mathbf{x})$ is a pre-defined perturbation set, to approximate the minimal empirical loss even under the strongest attack, directly learning the concept of robustness (Madry et al., 2018). Thus, the current AT objective can be formulated as a min-max problem :
@@
where $h_{\bm{\theta}}^{\text{DuRM}}(\mathbf{x}_{i})\in\mathbb{R}^{C_{d}}$ denotes the output logits of DuRM and $\mathbf{0}$ is a zero vector with the same dimension, which means there is no supervision information for these additional dummy classes.
@@ 2.2.2 Paradigm Formulation
where $\dot{\mathbf{y}}_{i}+\ddot{\mathbf{y}}_{i}$ equals to the original one-hot label vector $\mathbf{y}_{i}$. Then with $\mathbf{x}^{\prime}$ being the adversarial sample generated from $\mathbf{x}$ and $h_{\bm{\theta}}^{\text{Dummy}}(\mathbf{x}_{i})\in\mathbb{R}^{C}$ denoting the output logits from the $C$ appended dummy classes, our new AT paradigm can be formulated as:
@@
where $\dot{\mathbf{y}}_{i}^{\prime}$ and $\ddot{\mathbf{y}}_{i}^{\prime}$ can be differently weighted with $\dot{\mathbf{y}}_{i}$ and $\ddot{\mathbf{y}}_{i}$. For example, it would be valid to assign that $\dot{\mathbf{y}}_{i}^{(k)}\!+\ddot{\mathbf{y}}_{i}^{(k)}=0.5+0.5=1$ while $\dot{\mathbf{y}}_{i}^{\prime(k)}\!+\ddot{\mathbf{y}}_{i}^{\prime(k)}=1+0=1$.
At inference time, we can acquire the final output of the robust DNN classifier by projecting each dummy class to the corresponding original ones. As such a projection is separated from the computation graph of DNN and can be merely implemented in the run time, the robust mechanism can be simply explained as detecting (possible) adversarial attacks when certain samples are classified to the dummy classes and recovering such samples back to the corresponding original classes. We defer detailed discussion on the real-world threat model including adversary capacity to Section 3.1. Formally, with $\hat{y}$ denoting the final predicted class label, the projection can be formulated as:
$\hat{y}=\begin{cases}g_{\bm{\theta}}(\mathbf{x}_{i}),\!&g_{\bm{\theta}}(\mathbf{x}_{i})\!\leq C\\ g_{\bm{\theta}}(\mathbf{x}_{i})\!-C,\!&g_{\bm{\theta}}(\mathbf{x}_{i})\!>C\\ \end{cases}\hskip 5.0pt\text{where}\hskip 5.0ptg_{\bm{\theta}}(\mathbf{x}_{i})=\operatorname*{arg\,max}_{k=1,...,2\cdot C}\big{(}h_{\bm{\theta}}(\mathbf{x}_{i})\,\|\,h_{\bm{\theta}}^{\text{Dummy}}\!(\mathbf{x}_{i})\big{)}^{(k)}.$ (6)
@@
  • Approach. DuRM tends to adopt a smaller $C_{d}$ (e.g., typically 1 or 2) than $C$, and expects there should be no samples actually classified to the dummy classes. In contrast, our paradigm always adopts another $C$ dummy classes, clearly encouraging adversarial samples to be classified into them, and utilizing the one-to-one correspondences between original and dummy classes to detect and recover adversarial samples at inference time, thus acquiring certain robustness.
  • Achievement. Despite AT has been mentioned as one of several application scenarios of DuRM, as it actually does not involve any special designs for AT, the reported improvement is trivial (i.e., no significant advancements compared with baselines like PGD-AT). On the contrary, our approach achieves SOTA performance as shown in Section 3.
2.2.3 A Toy Case
Figure 4: A toy case using hard labels achieves fake robustness.
Before specifically proposing our AT method under the new paradigm formulated above, we first demonstrate a generalization problem crossing different adversaries through a toy case in this section, which also double-confirms the reasonability of introducing the soft label in our new paradigm. Specifically, we directly assign that $\dot{\mathbf{y}}_{i}=\ddot{\mathbf{y}}_{i}^{\prime}=\mathbf{y}$ and $\ddot{\mathbf{y}}_{i}=\dot{\mathbf{y}}_{i}^{\prime}=\mathbf{0}$, which means there is no supervision information to explicitly bridge corresponding benign and adversarial samples, and they are viewed to belong to absolutely different classes. We implement such a toy case with the same settings as the proof-of-concept experiments in Section 2.1, except additionally introducing Auto-Attack (Croce & Hein, 2020), which is one of the most solid adaptive attack ensembles for reliable test of robust generalization, as a more powerful adversary. As shown in Figure 4, our result demonstrates that although this toy case can achieve surprisingly high accuracy both on clean data and under PGD-10 adversary, it immediately collapses once facing Auto-Attack, which reveals that it fails to learn the real robustness that can generalize to unseen adversaries. On the contrary, it simply overfits the PGD-10 training samples. This toy case suggests that the most straightforward idea to completely separate the learning of benign and adversarial samples may be undesirable, implying the importance of building generalized correspondence between them. This not only supports the design of soft labels in our paradigm but also provides empirical guidance for our specific AT method proposed in the following Section 2.3.
@@ 2.3 Proposed Method: DUmmy Classes-based Adversarial Training
In this section, following our new AT paradigm, we specifically propose a novel AT method named DUmmy Classes-based Adversarial Training (DUCAT). Formally, for the target model $\bm{\theta}$ and certain benign sample $\mathbf{x}_{i}$, given $\mathbf{\hat{x}}_{i}^{\prime}$ representing the adversarial sample generated through $\mathbf{\hat{x}}_{i}^{\prime}=$ $\operatorname*{arg\,max}_{\,\mathbf{x}_{i}^{\prime}\in\mathcal{P}(\mathbf{x}_{i})}\mathds{1}(\operatorname*{arg\,max}_{k=1,...,C}h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})^{(k)}\neq y_{i})$, then based on Equations (5) and (6), our new empirical risk to be minimized can be formulated with the 0-1 loss (Zhang et al., 2019) as follows:
$\mathcal{R}^{\text{DUCAT}}\hskip 1.00006pt(\bm{\theta},\mathbf{x}_{i}):=$ $\alpha\cdot\big{(}\beta_{1}\cdot\mathds{1}(\,g_{\bm{\theta}}(\mathbf{x}_{i})\neq y_{i}\,)+(1\!-\!\beta_{1})\cdot\mathds{1}(\,g_{\bm{\theta}}(\mathbf{x}_{i})\neq(y_{i}\!+\!C))\big{)}$ (7)
@@
where $\mathds{1}(\cdot)$ represents the indicator function and the weights $\alpha$, $\beta_{1}$, $\beta_{2}$ are hyper-parameters.
@@
Such a deconstruction of our target risk is expected to facilitate the understanding of its principle and novelty. Specifically, existing AT methods either only focus on the adversarial risk $\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})\neq y_{i})$ or jointly considering the benign ones like $\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i})\neq y_{i})+\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})\neq y_{i})$ (Bai et al., 2021). Our toy case in Section 2.2.3 attributes benign and adversarial samples to completely irrelevant risks as $\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i})\neq y_{i})+\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})\neq(y_{i}\!+\!C))$, which is also confirmed undesirable. In contrast, our proposed risk expects the benign and adversarial samples with the same original label to be classified as an original class and a dummy one. The important difference is that neither the original class is necessary to be completely benign nor the dummy one is expected totally adversarial. The simple philosophy behind this idea is that assuming corresponding benign and adversarial samples to have completely different distributions is as unadvisable as assuming the same distributions between them. Instead, we utilize the weights $\beta_{1}$ and $\beta_{2}$ to inject our preference in a softer manner, suggesting that when learning the primary target label of a benign sample $\mathbf{x}_{i}$ (or adversarial sample $\mathbf{x}_{i}^{\prime}$) is found too hard for DNNs, automatically shunting them to the corresponding dummy (or benign) class is a more acceptable alternative option compared with other completely irrelevant classes. This approach is expected to not only reduce the overfitting to specific adversaries caused by the memorization of hard-label adversarial samples in the training process (Stutz et al., 2020; Dong et al., 2022; Cheng et al., 2022), but also establish explicit connections between corresponding benign and adversarial samples, laying a solid foundation for the inference time projection from dummy classes to the corresponding original ones as provided in Equation (6).
Table 1: The results of integrating the proposed DUCAT to four AT benchmarks on CIFAR-10, CIFAR-100 and Tiny-ImageNet with ResNet-18 under $\ell_{\infty}$ adversaries, clearly breaking the inherent trade-off between clean accuracy and robustness. All results are acquired from three runs.
Dataset Method Original +DUCAT (Ours)
   Clean PGD-10 PGD-100 Auto-Attack    Clean PGD-10 PGD-100 Auto-Attack
PGD-AT 82.92 51.81 50.34 46.74 88.81 65.10 62.71 58.61
CIFAR-10 TRADES 79.67 52.14 51.88 47.62 77.74 52.66 51.98 48.17
MART 77.93 53.61 52.83 46.70 80.65 58.42 57.81 50.18
Consistency-AT 83.42 53.20 51.68 47.72 89.51 66.83 63.80 57.18
PGD-AT 56.56 29.27 28.71 25.02 70.71 33.17 29.56 25.20
CIFAR-100 TRADES 55.39 29.61 29.28 24.51 55.41 30.69 30.38 25.23
MART 49.83 30.60 30.31 25.00 56.73 41.78 34.32 27.44
Consistency-AT 58.53 29.99 29.13 25.39 72.29 33.98 30.73 25.66
PGD-AT 46.32 21.75 21.52 17.07 56.18 24.23 22.54 18.68
Tiny-ImageNet TRADES 46.75 21.62 21.52 16.60 46.90 22.38 22.08 17.27
MART 39.70 22.98 22.79 17.18 43.37 25.23 25.68 18.41
Consistency-AT 46.54 22.58 21.70 17.60 59.64 25.40 23.22 18.92
@@
Finally, while 0-1 loss benefits conceptual analysis from the perspective of the risk, the optimization directly over it can be computationally intractable. So to end up with a real-world practical AT method, we introduce CE, the most commonly used for both supervised classification and conventional AT, as the surrogate loss to optimize the proposed risk $\mathcal{R}^{\text{DUCAT}}$ in Equation (7). It is also feasible to adopt more advanced loss functions such as boosted CE (Wang et al., 2020b), which creates conditions for integrating the proposed DUCAT with conventional AT methods. Formally, provided $l\,(\mathbf{y}_{i},\beta)$ being the non-one-hot soft label constructed as the supervision signal, such that:
@@
The overall objective of the proposed DUCAT can be formulated as:
@@
Then with CE $\mathcal{L}^{\text{CE}}$ serving as the surrogate loss $\mathcal{L}^{\text{surro}}$, we have:
@@
$\,\,\,\,\,-\!(1\!-\!\alpha)\!\cdot\!\big{(}\beta_{2}\!\cdot\!\log(\mathbf{p}^{(y_{i}\!+\!C)}\!(\mathbf{\hat{x}}_{i}^{\prime},\bm{\theta}))\!+\!(1\!-\!\beta_{2})\!\cdot\!\log(\mathbf{p}^{(y_{i})}\!(\mathbf{\hat{x}}_{i}^{\prime},\bm{\theta}))\big{)}\!\Big{)}\!,$
in which with $\sigma:\mathbb{R}^{2C}\!\rightarrow\!(0,1)^{2C}$ representing softmax, $\mathbf{p}(\mathbf{x}_{i},\bm{\theta})$ denotes the probabilistic prediction vector from DUCAT (i.e., with length $2C$ and the components adding up to one) such that:

### v2

@@ 2.2.1 Preliminaries
In clean training, for a classification task with $C\geq 2$ as the number of classes, given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1,...,n}$ with $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{1,...,C\}$ respectively denoting a natural sample and its supervised label, as well as $\mathbf{y}_{i}\in\mathbb{N}^{C}$ being the one-hot format of $y_{i}$, ERM aims to learn a classifier $h_{\bm{\theta}}:\mathbf{x}\rightarrow\mathbb{R}^{C}$, such that $h_{\bm{\theta}}(\mathbf{x}_{i})=\mathbf{q}(\mathbf{x}_{i},\bm{\theta})$ represents the output logits with each element $\mathbf{q}^{(k)}(\mathbf{x}_{i},\bm{\theta})$ corresponding to class $k$, to minimize the empirical risk $\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{D}}\,[\mathcal{L}(h_{\bm{\theta}}(\mathbf{x}),\mathbf{y})]$ through a classification loss $\mathcal{L}$ like cross-entropy (CE). Such a learning objective is formulated as:
@@
AT augments ERM to $\mathbb{E}_{\mathbf{x}^{\prime},\mathbf{y}\sim\mathcal{D}}\,[\max_{\,\mathbf{x}^{\prime}\in\mathcal{P}(\mathbf{x})}\mathcal{L}(h_{\bm{\theta}}(\mathbf{x}^{\prime}),\mathbf{y})]$, where $\mathbf{x}^{\prime}$ denotes the adversarial sample w.r.t. $\mathbf{x}$ and $\mathcal{P}(\mathbf{x})$ is a pre-defined perturbation set, to approximate the minimal empirical loss even under the strongest attack, directly learning the concept of robustness (Madry et al., 2018). Thus, the current AT objective can be reformulated as a min-max problem:
@@
where $h_{\bm{\theta}}^{\text{DuRM}}(\mathbf{x}_{i})\in\mathbb{R}^{C_{d}}$ denotes the output logits of DuRM and $\mathbf{0}$ is a zero vector in the same dimension, identifying no supervision information for these additional dummy classes. As discussed in Appendix G.1, although inheriting the term “dummy”, our work is different in essence from DuRM regarding motivation, significance and achievement. Especially, DuRM expects no samples classified to dummy classes, while ours is just the opposite as to be seen in Section 2.2.2. Thus, the specific significances behind “dummy” are completely different between DuRM and ours.
@@ 2.2.2 Paradigm Formulation
where $\dot{\mathbf{y}}_{i}+\ddot{\mathbf{y}}_{i}$ equals to the original one-hot label vector $\mathbf{y}_{i}$. With $\mathbf{x}^{\prime}$ being adversarial sample from $\mathbf{x}$ and $h_{\bm{\theta}}^{\text{Dummy}}(\mathbf{x}_{i})\in\mathbb{R}^{C}$ denoting the output logits from the $C$ appended dummy classes, our new AT paradigm can be formulated as:
@@
where $\dot{\mathbf{y}}_{i}^{\prime}$ and $\ddot{\mathbf{y}}_{i}^{\prime}$ can be differently weighted with $\dot{\mathbf{y}}_{i}$ and $\ddot{\mathbf{y}}_{i}$. For example, it would be valid to assign $\dot{\mathbf{y}}_{i}^{(k)}$ and $\ddot{\mathbf{y}}_{i}^{(k)}$ as 0.5, while assign $\dot{\mathbf{y}}_{i}^{\prime(k)}$ and $\ddot{\mathbf{y}}_{i}^{\prime(k)}$ respectively as 1 and 0, both of which add up to 1.
At inference time, we acquire the final output of the robust DNN classifier by projecting each dummy class to the corresponding original ones. As such a projection is separated from the computation graph of DNN and can be merely implemented in the run time, the robust mechanism can be simply explained as detecting (possible) adversarial attacks when certain samples are classified to the dummy classes and recovering such samples back to the corresponding original classes. We defer detailed discussion on the threat model including adversary capacity to Appendix C.2. With $\hat{y}$ denoting the final predicted class label, the projection is formulated as:
$\hat{y}=\begin{cases}g_{\bm{\theta}}(\mathbf{x}_{i}),\!&g_{\bm{\theta}}(\mathbf{x}_{i})\!\leq C\\ g_{\bm{\theta}}(\mathbf{x}_{i})\!-C,\!&g_{\bm{\theta}}(\mathbf{x}_{i})\!>C\\ \end{cases},\hskip 5.0pt\text{where}\hskip 5.0ptg_{\bm{\theta}}(\mathbf{x}_{i})=\operatorname*{arg\,max}_{k=1,...,2\cdot C}\big{(}h_{\bm{\theta}}(\mathbf{x}_{i})\,\|\,h_{\bm{\theta}}^{\text{Dummy}}(\mathbf{x}_{i})\big{)}^{(k)}.$ (6)
@@
2.2.3 Two-Hot Soft Label Matters
Within the new AT paradigm, apart from the dummy class, another novel concept proposed is the two-hot soft label. In this section, we demonstrate that abusing the new AT paradigm with hard labels rather than two-hot soft labels would lead to a generalization problem from training to test adversaries, which confirms the necessity and contribution of this design. Specifically, here we directly assign that $\dot{\mathbf{y}}_{i}=\ddot{\mathbf{y}}_{i}^{\prime}=\mathbf{y}$ and $\ddot{\mathbf{y}}_{i}=\dot{\mathbf{y}}_{i}^{\prime}=\mathbf{0}$, which means there is no supervision information to explicitly bridge corresponding benign and adversarial samples, and they are viewed to belong to absolutely different classes. We implement such a toy case with the same settings as the proof-of-concept experiments in Section 2.1, except additionally introducing Auto-Attack (AA) (Croce and Hein, 2020b) to evaluate
Figure 6: A toy case using hard labels achieves brittle robustness. The result is surprisingly high under training adversary PGD-10 but collapses under unseen AA.
the robust generalization. As shown in Figure 6, our result demonstrates that although this toy case can achieve surprisingly high accuracy both on clean data and under PGD-10 adversary, it immediately collapses once facing AA, which reveals that it fails to learn the real robustness that can generalize to unseen adversaries. On the contrary, it simply overfits the PGD-10 training samples. This toy case suggests that the most straightforward idea to completely separate the learning of benign and adversarial samples may be undesirable, implying the importance of building generalized correspondence between them. This not only supports the design of two-hot soft labels in our paradigm but also provides empirical guidance for our specific AT method proposed in the following Section 2.3. Besides, please kindly note that our two-hot soft labels are significantly different from the conventional soft labels, such as the ones in Müller et al. (2019), Shafahi et al. (2019), and Wu et al. (2024). We defer the detailed discussion to Appendix G.2.
@@ 2.3 Proposed Method: Dummy Classes-based Adversarial Training
Following our new AT paradigm, we specifically propose a novel AT method named DUmmy Classes-based Adversarial Training (DUCAT). Formally, for the target model $\bm{\theta}$ and certain benign sample $\mathbf{x}_{i}$, given $\mathbf{\hat{x}}_{i}^{\prime}$ representing the adversarial sample generated through $\mathbf{\hat{x}}_{i}^{\prime}=$ $\operatorname*{arg\,max}_{\,\mathbf{x}_{i}^{\prime}\in\mathcal{P}(\mathbf{x}_{i})}$ $\mathds{1}(\operatorname*{arg\,max}_{k=1,...,C}h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})^{(k)}\neq y_{i})$, then based on Equations (5) and (6), our new empirical risk to be minimized can be formulated with 0-1 loss (Zhang et al., 2019) and indicator function $\mathds{1}(\cdot)$ as:
$\mathcal{R}^{\text{DUCAT}}(\bm{\theta},\mathbf{x}_{i}):=$ $\alpha\cdot\big{(}\beta_{1}\cdot\mathds{1}(g_{\bm{\theta}}(\mathbf{x}_{i})\neq y_{i})+(1-\beta_{1})\cdot\mathds{1}(g_{\bm{\theta}}(\mathbf{x}_{i})\neq(y_{i}+C))\big{)}$ (7)
@@
where $\alpha$, $\beta_{1}$, $\beta_{2}$ are hyper-parameters, with discussion on their specific values deferred to Appendix F.
@@
Such a deconstruction of our target risk is expected to facilitate the understanding of its principle and novelty. Specifically, existing AT methods either only focus on the adversarial risk $\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})\neq y_{i})$ or jointly considering the benign ones like $\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i})\neq y_{i})+\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})\neq y_{i})$ (Bai et al., 2021). Our toy case in Appendix 2.2.3 attributes benign and adversarial samples to completely irrelevant risks as $\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i})\neq y_{i})+\mathds{1}(\,h_{\bm{\theta}}(\mathbf{x}_{i}^{\prime})\neq(y_{i}\!+\!C))$, which is also confirmed undesirable. In contrast, our proposed risk expects the benign and adversarial samples with the same original label to be classified as an original class and a dummy one. The important difference is that neither the original class is necessary to be completely benign nor the dummy one is expected totally adversarial.
The simple principle behind this idea is that assuming corresponding benign and adversarial samples to have completely different distributions is as unadvisable as assuming the same distributions between them. Instead, we utilize the weights $\beta_{1}$ and $\beta_{2}$ to inject our preference in a softer manner, suggesting that when learning the primary target label of a benign sample $\mathbf{x}_{i}$ (or adversarial sample $\mathbf{x}_{i}^{\prime}$) is found too hard for DNNs, automatically shunting them to the corresponding dummy (or benign) class is a more acceptable alternative option compared with other completely irrelevant classes. This approach is expected to not only reduce the overfitting to specific adversaries caused by the memorization of hard-label training adversarial samples (Stutz et al., 2020; Dong et al., 2022; Cheng et al., 2022), but also establish explicit connections between corresponding benign and adversarial samples, laying a solid foundation for the inference time projection from dummy classes to the corresponding original ones as provided in Equation (6).
@@
Finally, while 0-1 loss benefits conceptual analysis from the perspective of the risk, the optimization directly over it can be computationally intractable. So to end up with a real-world practical AT method, we introduce CE, the most commonly used for both supervised classification and conventional AT, as the surrogate loss to optimize the proposed risk $\mathcal{R}^{\text{DUCAT}}$ in Equation (7). It is also feasible to adopt more advanced loss functions such as boosted CE (Wang et al., 2020b), which creates conditions for integrating the proposed DUCAT with conventional AT methods. Formally, provided that $l\,(\mathbf{y}_{i},\beta)$ denotes the two-hot soft label constructed as the supervision signal, such that:
@@
The overall objective of DUCAT can be formulated as:
@@
Then with $\mathcal{L}^{\text{CE}}$ serving as the surrogate loss $\mathcal{L}^{\text{surro}}$, we have:
@@
$\hskip 0.90005pt-(1-\alpha)\cdot\big{(}\beta_{2}\cdot\log(\mathbf{p}^{(y_{i}+C)}(\mathbf{\hat{x}}_{i}^{\prime},\bm{\theta}))+(1-\beta_{2})\cdot\log(\mathbf{p}^{(y_{i})}(\mathbf{\hat{x}}_{i}^{\prime},\bm{\theta}))\big{)}\Big{)},$
in which with $\sigma:\mathbb{R}^{2C}\!\rightarrow\!(0,1)^{2C}$ representing the softmax, $\mathbf{p}(\mathbf{x}_{i},\bm{\theta})$ denotes the output probabilistic prediction from DUCAT (i.e., with the length of $2C$ and the components adding up to one), such that:
@@
Table 1: The results of integrating the proposed DUCAT to four AT benchmarks on CIFAR-10, CIFAR-100 and Tiny-ImageNet with ResNet-18 under $\ell_{\infty}$ adversaries, clearly releasing the inherent trade-off between clean accuracy and robustness. All results are acquired from three runs.
Dataset Method Clean PGD-10 PGD-100 AA Mean NRR
PGD-AT 82.92 51.81 50.34 46.74 64.830 59.782
      + DUCAT 88.81 (+7.1%) 65.10 (+25.7%) 62.71 (+24.6%) 58.61 (+25.4%) 73.710 (+13.7%) 70.617 (+18.1%)
TRADES 79.67 52.14 51.88 47.62 63.645 59.610
      + DUCAT 77.74 (-2.4%) 52.66 (+1.0%) 51.98 (+0.2%) 48.17 (+1.2%) 62.955 (-1.1%) 59.483 (-0.2%)
CIFAR-10 MART 77.93 53.61 52.83 46.70 62.315 58.402
      + DUCAT 80.65 (+3.5%) 58.42 (+9.0%) 57.81 (+9.4%) 50.18 (+7.5%) 65.415 (+5.0%) 61.867 (+5.9%)
Cons-AT 83.42 53.20 51.68 47.72 65.570 60.711
      + DUCAT 89.51 (+7.3%) 66.83 (+25.6%) 63.80 (+23.5%) 57.18 (+19.8%) 73.345 (+11.8%) 69.782 (+14.9%)
Average 80.99 52.69 51.68 47.20 64.090 59.626
      + DUCAT 84.18 (+3.9%) 60.75 (+15.3%) 59.08 (+14.3%) 53.54 (+13.4%) 68.856 (+7.4%) 65.437 (+9.7%)
PGD-AT 56.56 29.27 28.71 25.02 40.790 34.693
      + DUCAT 70.71 (+25.0%) 33.17 (+13.3%) 29.56 (+3.0%) 25.20 (+0.7%) 47.955 (+17.6%) 37.158 (+7.1%)
TRADES 55.39 29.61 29.28 24.51 39.950 33.983
      + DUCAT 55.41 (+0.0%) 30.69 (+3.6%) 30.38 (+3.8%) 25.23 (+2.9%) 40.320 (+0.9%) 34.672 (+2.0%)
CIFAR-100 MART 49.83 30.60 30.31 25.00 37.415 33.295
      + DUCAT 56.73 (+13.8%) 41.78 (+36.5%) 34.32 (+13.2%) 27.44 (+9.8%) 42.085 (+12.5%) 36.989 (+11.1%)
Cons-AT 58.53 29.99 29.13 25.39 41.960 35.417
      + DUCAT 72.29 (+23.5%) 33.98 (+13.3%) 30.73 (+5.5%) 25.66 (+1.1%) 48.975 (+16.7%) 37.876 (+6.9%)
Average 55.08 29.87 29.36 24.98 40.029 34.347
      + DUCAT 63.79 (+15.8%) 34.91 (+16.9%) 31.25 (+6.4%) 25.88 (+3.6%) 44.834 (+12.0%) 36.674 (+6.8%)
PGD-AT 46.32 21.75 21.52 17.07 31.695 24.947
      + DUCAT 56.18 (+21.3%) 24.23 (+11.4%) 22.54 (+4.7%) 18.68 (+9.4%) 37.430 (+18.1%) 28.037 (+12.4%)
TRADES 46.75 21.62 21.52 16.60 31.675 24.500
      + DUCAT 46.90 (+0.3%) 22.38 (+3.5%) 22.08 (+2.6%) 17.27 (+4.0%) 32.085 (+1.3%) 25.244 (+3.0%)
Tiny-ImageNet MART 39.70 22.98 22.79 17.18 28.440 23.982
      + DUCAT 43.37 (+9.2%) 25.23 (+9.8%) 25.68 (+12.7%) 18.41 (+7.2%) 30.890 (+8.6%) 25.848 (+7.8%)
Cons-AT 46.54 22.58 21.70 17.60 32.070 25.541
      + DUCAT 59.64 (+28.1%) 25.40 (+12.5%) 23.22 (+7.0%) 18.92 (+7.5%) 39.280 (+22.5%) 28.727 (+12.5%)
Average 44.83 22.23 21.88 17.11 30.970 24.743
      + DUCAT 51.52 (+14.9%) 24.31 (+9.3%) 23.38 (+6.8%) 18.32 (+7.1%) 34.921 (+12.8%) 26.964 (+9.0%)